OpenStudy (anonymous):

what is the x-intercept for the equation -2x+5y=-10

OpenStudy (anonymous):

set y equal to zero and solve for x -2x + 5y = -10 -2x = -10 divide both sides by -2 x = 5

OpenStudy (anonymous):

i think

OpenStudy (kamibug):

That's correct. :) \(\Huge\bf\color{seagreen}{Welcome~to~Openstudy!}\)